Write a system of equations. Solve the system.
the two top-grossing movies of 2003 were the lord of the rings and finding nemo. Finding nemo grossed $21.4 million less than The lord of the rings, and together the two films took in $700.8 million. How much did each of these movies earn? Show work.

Answer by oberobic(2304)

L = revenue earned by the Lord of the Rings
F = revenue earned by Finding Nemo
Given:
L + F = 700.8
L = F + 21.4
Subtract F from both sides
L - F = 21.4
.
Now we have to equations:
L + F = 700.8
L - F = 21.4
.
Add them
2L = 722.2
Divide both sides by 2
L = 361.1
.
Looking back, we can substitute L = 361.1 to find F.
L - F = 21.4
361.1 -F = 21.4
Add F to both sides
361.1 = F +21.4
Subtract 21.4 from both sides
361.1 - 21.4 = F, which means
F = 361.1 -21.4
F = 339.7
.
Always check your work!
.
In this case use the other equation you were given.
L + F = ??
361.1 + 339.7 = 700.8
That's right.
.
Answer: The Lord of the Rings earned $361.1 million; Finding Nemo earned $339.7 million.
.
Done.